OpenStudy (anonymous):

The local zoo has two water tanks for the elephant exhibit that are leaking One water tank contains 12 gal of water and is leaking at a constant rate of 3 gal/h. The second water tank contains 8 gal of water and is leaking at a constant rate of 5 gal/h. When will the two tanks have the same amount of water? Explain. Let x = the number of hours the tanks are filling and let y = the number of gallons in the tank. A. In 2 hours, because the solution to the system is (2,18). B. They will never have the same amount of water because the solution to the system is (–2,18). It is not possible to

OpenStudy (anonymous):

A. In 2 hours, because the solution to the system is (2,18). B. They will never have the same amount of water because the solution to the system is (–2,18). It is not possible to have time be –2 hours. C. In –2 hours, because the solution to the system is (–2,18). D. They will never have the same amount of water because the solution to the system is (–2,18). It is not possible to have –2 gallons in the tanks.

OpenStudy (anonymous):

ill give a MEDAL to anyone who can help!!!

OpenStudy (imstuck):

The way you solve this is in this way. You have your first equation that looks like this: y = 12 - 3x And then you have the second equation that looks like this: y = 8 - 5x x is the number of hours and y is the amount of water.

OpenStudy (imstuck):

since you are wanting to know when the tanks will have the same amount of water, y, you set the equations equal to one another. That means that the y, the amount of water in both equations, is the same, which is what you're looking for...when they are the same.

OpenStudy (imstuck):

12-3x=8-5x 4 = -2x -2 = x. Now remember that x is the number of hours it will take for the tanks to have an equal amount of water. Since time cannot be measured in the negative, the tanks will never have the same amount of water.It is B from your choices.

OpenStudy (anonymous):

thank you so much that actually makes sense to me

OpenStudy (imstuck):

Ok, you're welcome. medal me? ; )

OpenStudy (anonymous):

of course =] just have to figure out how im on my phone and it looks different.

OpenStudy (anonymous):

found it. it was hiding